# ALERT Open Dataset and Input-Size-Agnostic Vision Transformer for Driver Activity Recognition using IR-UWB

### v1

@@ III-E Dataset Comparison
In UWB DAR, the existence of standard open dataset is significant to accelerate the studies. In [RaDA], they provide the first open dataset, RaDA, for UWB DAR including comprehensive 6 activities such as driving, autopilot, sleeping, driving & smartphone utilization, smartphone utilization, and talking to a passenger. However, it is not enough to serve as de facto standard open dataset compared to our dataset, ALERT. The detailed differences between our dataset and RaDA are as follows and summarized in Table IV.
  • Activities: We select 6 distracted driving activities including Smoke, Drink, and Panel, which are observed more frequently and cause more accidents than the 5 activities identified in RaDA.
@@
    Features ALERT RaDA [RaDA]
@@
  • Restrictions of activities: As mentioned above, we provide only minimal guidance to volunteers to prevent intentionally unusual behavior. However, in RaDA, they ask for some restrictions on activities. For instance, volunteers should use their right hand, turn their heads right when talking to a passenger, and use both hands when operating a smartphone. Such instructions may make the dataset more uniform and easier to train with, but they do not reflect typical real-world behavior.
  • Observation window size of dataset: RaDA’s target activities can be recognized within a second to deal with their activities. However, we recommend a sufficient observation window size, at least 5 seconds, to identify more complex activities in the ALERT dataset; we demonstrate this in Section VI. A sufficient observation period is crucial to accurately capturing such activities’ characteristics in DAR scenario. Consequently, although both datasets provide comparable samples, the ALERT offers a longer total observation time.
  • Data collection environment: In RaDA, the dashboard props up the radar installed in front of the driver, obstructing the driver’s view and leading to severe accidents. However, this setup is possible because the RaDA dataset has been collected in a simulated environment. In contrast, we position the radar on the air vent to avoid visual obstruction and collect our dataset in a real-driving environment. This setting incorporates more realistic factors, such as ground and engine vibrations, and RF signal interference commonly encountered in urban areas. Consequently, the ALERT dataset offers a more realistic representation than RaDA.
  • Comprehensive data: In RaDA, they only provide the single domain data. However, our ALERT dataset offers both range and frequency domain data, enabling the use of heterogeneous domains. Utilizing both domains contributes to extensive benchmarking experiments and evaluations. Additionally, our dataset allows manipulation according to user demands, including freely adjusting the cropping area of range and frequency and observation window size. Consequently, we can conduct extensive and varied experiments using the ALERT dataset, with the results evaluated and presented in Section VI.
@@ III-F Dataset and Code Availability
We publicly release the ALERT dataset on Figshare to facilitate further research and experimentation. Additionally, we provide the benchmarking code on GitHub, enabling users to verify benchmarking performance and offering guidance on utilizing the code with the ALERT dataset. The URLs for the dataset and benchmarking codes are as follows,
ALERT dataset: https://doi.org/10.6084/m9.figshare.28244525.v3
Benchmarking codes: https://github.com/ALERTdataset/ALERT.git
@@ IV Benchmkaring Design
To provide benchmarking performance, we design various learning models including CNN-based [googlenet, resnet, densenet, mobilenet], RNN-based [maitre2020fall, maitre2021recognizing, noori2021ultra], and transformer-based approaches [deit, vit]. Additionally, we incorporate a multi-view learning (MVL)-based domain fusion method into all designs to demonstrate the scalability of using the ALERT dataset and to achieve better performance in DAR.
@@ IV-A Domain Fusion
In IR-UWB radar sensing, we primarily obtain range-time information between an object and the radar. However, relying solely on range information makes it challenging to extract distinguishing features needed to recognize similar behaviors. For example, smoking and drinking while driving are very similar activities: to do these, the driver raises and lowers their hand in a similar motion. Similarly, using a phone while driving can make it difficult to distinguish safe driving behavior. These actions exhibit only very subtle differences in range domain features. In DAR, which involves comprehensive and complex activities, failing to accurately classify similar activities can lead to decreased overall performance.
To distinguish similar activities, we implement a domain fusion method that utilizes both range and frequency domain data to capture complementary features obtained from UWB sensing. For effective fusion of range and frequency domains, it is essential to understand the feature characteristics of each domain. Each domain has distinct characteristics, outlined as follows.
The range-time domain data captures spatial information by mapping the distance of objects or activities from the sensor over time. This data is highly effective for tracking precise movements, especially rapid actions, as it can detect small shifts in position. Additionally, range-time data includes reflections from walls or objects, which can provide extra details about the surrounding environment.
We can obtain frequency-time information through a Fourier transform of the range-time data. It focuses on changes in Doppler shift over time, offering insights into the speed and direction of movement. While it provides less detail about position, frequency-time data is better suited for tracking movement velocity and patterns. Furthermore, frequency-time data is less affected by reflections, resulting in cleaner data that is easier to analyze for movement patterns. By integrating the two domains, we can make more accurate and reliable judgments about confusing labels.
In domain fusion, we leverage the various features obtained from multiple domains and concatenate each feature to classify the object. Similarly, some works [HARSANET, SleepPoseNet] apply the MVL-based domain fusion method to the UWB HAR for better performance. They transform the same UWB data into both range domain and frequency domain components, separately process each branch’s range and frequency inputs, and then concatenate to classify the activities. The details of the domain fusion design are provided in the following, and the evaluation results using both single-domain datasets and domain fusion datasets are presented in Section VI.
@@ IV-B Design for CNN-based Approach
Figure 6: Design of learning algorithms for benchmarking.
We benchmark the performance of various CNN architectures including GoogLeNet [googlenet], ResNet [resnet], DenseNet [densenet], and MobileNet [mobilenet]. The aforementioned algorithms are CNN-based feature extractors as depicted in Fig. 6. The range and frequency data are fed into the CNN-based feature extractor. After processing through the feature extractor, features for range and frequency are concatenated and then inputted into a classifier to recognize the activity. The classifier comprises three fully connected layers, with various dimensions tailored to the specific algorithm employed.
@@ IV-C Design for RNN-based Approach
To design the RNN-based approach, we follow the conceptual framework introduced in [maitre2020fall, maitre2021recognizing, noori2021ultra]. Although the ALERT dataset includes sufficiently long sequences for temporal modeling, we do not adopt xLSTM [xlstm] for benchmarking. This is because xLSTM processes individual UWB pulses sequentially, and each pulse contains only limited information. In UWB radar, meaningful activity patterns emerge only when multiple pulses are accumulated over time. To address this, we adopt the CNN+LSTM architecture that groups several pulses into short snippets, extracts spatial features from each snippet using a CNN, and then models the temporal dependencies between these snippets using an LSTM.
@@ IV-D Design for Transformer-based Approach
For a Transformer-based approach, we utilize Vision Transformer (ViT) model [vit] and Data-efficient Image Transformer (DeiT) model [deit], which is a variant of ViT.
In the transformer-based approach, we adopt the early fusion method, as illustrated in Fig. 6. Since transformer-based models are computationally intensive, applying the late fusion method in a real-time DAR system is challenging. Late fusion requires two independent ViT models for each input, followed by concatenating the features extracted from their outputs, which significantly increases computational overhead.
@@ V Input-Size-Agnostic Vision Transformer (ISA-ViT)
Transferring the remarkable capabilities of ViTs to the UWB domain is critical for effectively classifying 2D image-like UWB datasets. However, applying a ViT pre-trained on ImageNet [imagenet] to UWB data presents a significant challenge. In image processing, ViTs typically operate on an input size of $224\times 224$, divided into $16\times 16$ patches, resulting in 196 patches (14 along both the height and width of the image). Each patch is transformed into a patch embedding vector and combined with the positional embedding vector (PEV) containing relative spatial information, forming the input to the transformer encoder.
However, in practice, the input size of UWB data cannot always be guaranteed to match the fixed size of $224\times 224$. This discrepancy makes it infeasible to directly apply pre-trained PEVs to UWB data. To tackle this critical issue, we propose the ISA-ViT, a novel approach specifically designed to adapt ViTs for UWB data and introduce the details of ISA-ViT as follows.
@@ V-A Preliminary Studies
We can refer to [ast] to manipulate the PEVs, as mentioned above the second approach. They calculate the number of patches (using $\times16$ patches) along the height and width of the input data. If the number of patches in the side exceeds 14, they apply interpolation to the PEVs for height; otherwise, they use truncation. We implement manipulating PEVs by following the conceptual approach used in [ast].
@@
Method Dataset ALERT RaDA [RaDA]
@@
Through the preliminary studies, we observe the importance of preserving the pre-trained $14\times 14$ PEV sequence for non-image input sizes. In the following section, we introduce ISA-ViT, a detailed method for maintaining the pre-trained sequence of PEVs, designed to effectively apply PEVs to UWB data.
@@ V-C ISA-ViT
1:Parameters:
@@
[…]
11:Input: $data_{UWB}$
12:Initialize:
[…]
$16k_{original}\leftarrow$ 16
17:Calculation of $k$ and $side_{extended}$:
[…]
20:Resizing input:
[…]
24:Averaging channel and resizing kernel:
[…]
29:Linear projection:
[…]
31:Output: Patch embedding vectors
@@ V-D Domain Fusion of ISA-ViT
For domain fusion, we might initially consider using two parallel ViT models to equally extract features from range and frequency data. However, simply fusing the features from both ViT models would make the system quite heavy as mentioned in Section IV-D. To consider both accuracy and system responsiveness, we devise the light-weight late fusion method as illustrated in Fig. 8(b). We apply ISA-ViT to the range data and use a lightweight feature extractor model for the frequency data. The range data is divided into $k\times k$ patches, specifically $14\times 14$ patches, which are transformed into patch embedding vectors via linear projection. These vectors are then combined with pre-trained PEVs and processed through the transformer encoder to obtain the feature representation of the range domain data. In contrast, we pass the frequency data through the lightweight feature extractor without resizing, further reducing computational cost. This approach not only optimizes system efficiency but also places greater emphasis on the range domain data due to its distinctive feature characteristics.
The reason we focus more on range domain data is that it provides essential information about behaviors. When identifying a driver’s actions while driving, range information is highly valuable, as it helps in understanding the size and nature of behaviors. For example, by utilizing range information, we can distinguish whether the driver is holding a phone in front of their body to use it or bringing a cigarette to their mouth to smoke. While frequency information offers insights into speed changes and movement direction, it is insufficient on its own to accurately identify the driver’s behavior. For instance, even if the speed remains constant, frequency data alone cannot reveal whether the driver is dozing off or actively focused on driving. Consequently, we place greater emphasis on range domain features and less on frequency domain features.
We primarily utilize features extracted from the range domain, with frequency domain data providing supplementary information. Instead of concatenating frequency features at the same ratio as range features, we apply an adjusting factor $\beta$ to balance the frequency domain features before concatenation. The rationale for this approach is that range and frequency data do not always complement each other positively in domain fusion. This method ensures that the contribution from the frequency domain is balanced and does not overpower the more informative range domain features. We can determine the optimal value of the adjusting factor $\beta$ through the training phase.
@@ VI-A Experimental Settings
Using the ALERT dataset, we benchmark several learning algorithms, including GoogLeNet [googlenet], ResNet [resnet], DenseNet [densenet], MobileNet [mobilenet], an RNN-based model [maitre2020fall, maitre2021recognizing, noori2021ultra] (denoted as RNN), and transformer-based models such as DeiT [deit] and ViT [vit]. Each model is implemented as described in Section IV.
To ensure adequate fine-tuning, we train the models for 30 epochs, using a learning rate between 0.00001 and 0.0001 to achieve convergence of the training loss for each algorithm. Cross-entropy loss is employed as the loss function, and the Adam optimizer is used for all models.
Evaluation of ISA-ViT. We thoroughly evaluate the performance of DAR using ISA-ViT and the proposed domain fusion method. To conduct extensive experiments, we employ the ALERT dataset and RaDA dataset [RaDA] which are open dataset leveraging IR-UWB to detect driver distraction activities.
@@
For model training, we strictly separate the training and testing datasets; the testing dataset consists of data from drivers who were not seen during the training phase. We employ a learning rate as 0.00001 to ensure convergence the training loss. The base model is a ViT pre-trained on ImageNet [imagenet], which uses embedding vectors of size 1,024. We fine-tune the model over 30 epochs, optimizing it with cross-entropy loss and the Adam optimizer to ensure robust performance.
VI-B Benchmarking Results for ALERT Dataset
VI-B1 Benchmarking of learning algorithms
For rigorous evaluation, we conduct benchmark tests using an unseen dataset during the training phase. In these experiments, we fine-tune pre-trained models, leveraging the pre-trained models’ great capabilities for feature extraction effectively to the UWB domain.
To identify a suitable model for UWB DAR, we first present the average accuracy of each algorithm. We select four unseen testers and perform leave-one-out cross-validation to calculate their average accuracy. Fig. 9(a) shows the average accuracy of various learning algorithms.
CNN-based approaches achieve an accuracy range of 61.37–71.22%, outperforming both RNN-based and transformer-based approaches. The underperformance of the RNN-based approach stems from the lack of a pre-trained LSTM model. While it leverages pre-trained models for snippet feature extraction, the sequential LSTM model itself is not pre-trained, which limits its overall effectiveness.
Although transformer-based approaches are state-of-the-art models, they also fail to outperform CNN-based methods due to challenges outlined in Section V. While transformers excel in extracting features for image-based tasks, they require resizing UWB data and adapting pre-trained PEVs. Without addressing these issues, transformer-based approaches cannot fully leverage their potential in the UWB domain.
In contrast, the proposed ISA-ViT overcomes these challenges by resizing UWB data without information loss and adapting the pre-trained PEVs to the resized data. As a result, ISA-ViT achieves the highest accuracy of 76.28%, demonstrating its superior performance.
@@ VI-B1 Benchmarking of learning algorithms
TABLE VI: Benchmarking [precision / recall] of learning algorithms
@@
Along with the accuracy, other important metrics to consider are the precision and recall for each activity. High precision or recall does not always guarantee good accuracy, but they are valuable metrics for evaluating classification performance. Table VI presents the precision and recall values for each activity across the evaluated learning algorithms. The underlined values in Table VI represent the fluctuations in precision and recall for each activity. The variance of these metrics is particularly important as it reflects the model’s ability to maintain stable classification performance across different activities. Fig. 9(b) the mean and deviation of activities for each metric by algorithms. As shown in the figure, other algorithms often exhibit significant deviations; the ISA-ViT demonstrates a generally stable variance, ensuring more consistent classification performance for each activity.
VI-C Ablation Studies for ALERT dataset
We conduct various experiments to identify the optimal parameters, trends, and trade-offs for each algorithm when utilizing the ALERT dataset.
VI-C1 Analysis for manipulation of data size

### v2

@@ III-E Dataset Comparison
In UWB DAR, the existence of standard open dataset is significant to accelerate the studies. In [6], they provide the first open dataset, RaDA, for UWB DAR including comprehensive 6 activities such as driving, autopilot, sleeping, driving & smartphone utilization, smartphone utilization, and talking to a passenger. However, it is not enough to serve as de facto standard open dataset compared to our dataset, ALERT. The detailed differences between our dataset and RaDA are as follows and summarized in Table IV. Compared to RaDA, ALERT dataset is collected in real driving with an air-vent mounting that maintains a stable sensing geometry without interfering with the driver’s view. In addition, ALERT provides longer samples (5 s) and more fine-grained activity labels designed to reflect natural in-vehicle behaviors with fewer restrictive execution rules. Moreover, ALERT offers both range-time and frequency-time representations, enabling flexible input manipulations (e.g., cropping and window/size variations) for systematic evaluations. Finally, ALERT includes a broader set of driving-induced motion conditions (e.g., stop-and-go and road-surface variations), which helps assess robustness under realistic vibration and traffic dynamics.
@@
Features ALERT RaDA [6]
@@ III-F Dataset and Code Availability
We publicly release the ALERT dataset to support further research and experimentation. We also provide the dataset processing code on GitHub, along with guidelines for using it with the ALERT dataset. The URLs for the dataset and codes are as follows, ALERT dataset: https://github.com/ALERTdataset/ALERT.git
@@ IV Benchmkaring Design
Figure 6: Design of learning algorithms for benchmarking.
To provide benchmarking performance, we design various learning models including CNN-based [46, 17, 20, 18], RNN-based [33, 34, 38], and transformer-based approaches [47, 11]. Additionally, we incorporate a multi-view learning (MVL)-based domain fusion method into all designs to demonstrate the scalability of using the ALERT dataset and to achieve better performance in DAR.
@@ IV-A Domain Fusion
In IR-UWB sensing, range-time captures fine-grained spatial motion but often provides insufficient cues to separate visually similar driver actions (e.g., Smoke vs. Drink). Frequency-time, obtained via Fourier transform, emphasizes Doppler-driven velocity patterns and is typically less sensitive to multipath reflections, yielding cleaner signatures for motion dynamics. We therefore fuse both domains to leverage complementary cues and improve discrimination of confusing labels.
In domain fusion, we leverage the various features obtained from multiple domains and concatenate each feature to classify the object. Similarly, some works [9, 42] apply the MVL-based domain fusion method to the UWB HAR for better performance. They transform the same UWB data into both range domain and frequency domain components, separately process each branch’s range and frequency inputs, and then concatenate to classify the activities. The details of the domain fusion design are provided in the following, and the evaluation results using both single-domain datasets and domain fusion datasets are presented in Section VI.
@@ IV-B Design for CNN-based Approach
We benchmark the performance of various CNN architectures including GoogLeNet [46], ResNet [17], DenseNet [20], and MobileNet [18]. The aforementioned algorithms are CNN-based feature extractors as depicted in Fig. 6. The range and frequency data are fed into the CNN-based feature extractor. After processing through the feature extractor, features for range and frequency are concatenated and then inputted into a classifier to recognize the activity. The classifier comprises three fully connected layers, with various dimensions tailored to the specific algorithm employed.
@@ IV-C Design for RNN-based Approach
To design the RNN-based approach, we follow the conceptual framework introduced in [33, 34, 38]. Although the ALERT dataset includes sufficiently long sequences for temporal modeling, we do not adopt xLSTM [5] for benchmarking. This is because xLSTM processes individual UWB pulses sequentially, and each pulse contains only limited information. In UWB radar, meaningful activity patterns emerge only when multiple pulses are accumulated over time. To address this, we adopt the CNN+LSTM architecture that groups several pulses into short snippets, extracts spatial features from each snippet using a CNN, and then models the temporal dependencies between these snippets using an LSTM.
@@ IV-D Design for Transformer-based Approach
For a Transformer-based approach, we utilize Vision Transformer (ViT) model [11] and Data-efficient Image Transformer (DeiT) model [47], which is a variant of ViT. In the transformer-based approach, we adopt the early fusion method, as illustrated in Fig. 6. Since transformer-based models are computationally intensive, applying the late fusion method in a real-time DAR system is challenging. Late fusion requires two independent ViT models for each input, followed by concatenating the features extracted from their outputs, which significantly increases computational overhead.
@@ V Input-Size-Agnostic Vision Transformer (ISA-ViT)
Transferring the remarkable capabilities of ViTs to the UWB domain is critical for effectively classifying 2D image-like UWB datasets. However, applying a ViT pre-trained on ImageNet [10] to UWB data presents a significant challenge. Beyond the input-size mismatch, transferring image-pretrained ViTs to UWB radar also involves a domain gap (i.e., different signal statistics and physical meaning), which can limit “plug-and-play” transfer and motivates fine-tuning on target UWB data. In image processing, ViTs typically operate on an input size of $224\times 224$, divided into $16\times 16$ patches, resulting in 196 patches (14 along both the height and width of the image). Each patch is transformed into a patch embedding vector and combined with the positional embedding vector (PEV) containing relative spatial information, forming the input to the transformer encoder. This 14×14 token grid is an empirically established standard in pre-trained ViT model [11] and is tightly coupled with the learned positional embeddings, making its preservation important for stable transfer learning.
However, in practice, the input size of UWB data cannot always be guaranteed to match the fixed size of $224\times 224$. This discrepancy makes it infeasible to directly apply pre-trained PEVs to UWB data without altering the original token grid, which may degrade the effectiveness of transfer learning. Moreover, naïve up/down-sampling to fit 224×224 can discard signal details and weaken spatiotemporal structures, especially when the original input is much larger. To tackle this critical issue, we propose the ISA-ViT, a novel approach specifically designed to adapt ViTs for UWB data and introduce the details of ISA-ViT as follows.
@@ V-A Preliminary Studies
We can refer to [14] to manipulate the PEVs, as mentioned above the second approach. They calculate the number of patches (using $\times16$ patches) along the height and width of the input data. If the number of patches in the side exceeds 14, they apply interpolation to the PEVs for height; otherwise, they use truncation. We implement manipulating PEVs by following the conceptual approach used in [14].
@@
Method Dataset ALERT RaDA [6]
@@
Through the preliminary studies, we observe the importance of preserving the pre-trained $14\times 14$ PEV sequence for non-image input sizes. These results highlight that better performance is achieved when the pre-trained PEVs are reused in their original structure while the input is adapted in an information-preserving manner. This insight directly motivates our design goal in ISA-ViT: to enable effective utilization of pre-trained PEVs for highly asymmetric UWB inputs by minimizing resizing-induced information loss and maintaining the pre-trained PEV sequence. In the following section, we introduce ISA-ViT, a detailed method for maintaining the pre-trained sequence of PEVs, designed to effectively apply PEVs to UWB data.
@@ V-C ISA-ViT
1:Input: $data_{UWB}$
2:Output: Patch embedding vectors
3:Initialize: $num_{patches}\!\leftarrow\!14$, $k_{original}\!\leftarrow\!16$
4:Input size and patch-size calculation:
@@
13:Linear projection:
14:return $CNN\!\left(data_{UWB},\,weight_{kernel}(P_{k})\right)$
@@ V-D Domain Fusion of ISA-ViT
A straightforward approach for domain fusion is to use two parallel ViTs to extract features from the range and frequency domains. However, directly fusing two ViT backbones results in a heavy model, as discussed in Sec. IV-D. To balance accuracy and responsiveness, we propose a lightweight late-fusion scheme (Fig. 8(b)): we apply ISA-ViT to the range data, while using a lightweight feature extractor for the frequency data. The range input is partitioned into $k\times k$ patches (i.e., $14\times 14$), projected into patch embeddings, combined with pre-trained PEVs, and encoded by the transformer to obtain a range-domain representation. In contrast, the frequency input is fed into the lightweight extractor without resizing, which further reduces computation. This design improves efficiency while prioritizing the more informative range-domain cues.
We emphasize the range domain because it provides essential spatial context for behavior understanding. Range information helps distinguish actions with similar motion patterns but different spatial configurations (e.g., holding a phone in front of the body versus bringing a cigarette to the mouth). In contrast, frequency information mainly reflects motion dynamics (e.g., speed changes and movement direction) and is often insufficient alone; for example, with similar motion speeds, frequency data cannot reliably differentiate dozing off from attentive driving. Therefore, we treat frequency features as complementary cues and assign greater weight to range-domain features in fusion.
We primarily utilize features extracted from the range domain, with frequency domain data providing supplementary information. Instead of concatenating frequency features at the same ratio as range features, we apply an adjusting factor $\beta$ to balance the frequency domain features before concatenation. The contribution of frequency-domain features is not known a priori; therefore, we scale the frequency-domain feature by $\beta$ prior to concatenation:
$\mathbf{z}=[\mathbf{feat}_{range};\beta\times\mathbf{feat}_{freq.}],$ (4)
where $\mathbf{feat}_{range}$ and $\mathbf{feat}_{freq.}$ denote the range- and frequency-domain feature vectors, respectively. The rationale for this approach is that range and frequency data do not always complement each other positively in domain fusion. This method ensures that the contribution from the frequency domain is balanced and does not overpower the more informative range domain features. We treat $\beta$ as a trainable scalar parameter and optimize it jointly with the network parameters via backpropagation during training.
@@ VI-A Experimental Settings
Using the ALERT dataset, we benchmark several learning algorithms, including GoogLeNet [46], ResNet [17], DenseNet [20], MobileNet [18], an RNN-based model [33, 34, 38] (denoted as RNN), and transformer-based models such as DeiT [47] and ViT [11]. Each model is implemented as described in Section IV.
To ensure adequate fine-tuning, we train the models for 30 epochs, using a learning rate between 0.00001 and 0.0001 to achieve convergence of the training loss for each algorithm. Cross-entropy loss is employed as the loss function, and the Adam optimizer is used for all models. To ensure a fair comparison, we fix the train and test batch sizes to 25 for all models and set the adaptation batch size to 5. In addition, features extracted from each backbone are fed into the same classifier head for classification, and no weight decay, dropout, or additional regularization is applied.
Evaluation of ISA-ViT. We thoroughly evaluate the performance of DAR using ISA-ViT and the proposed domain fusion method. To conduct extensive experiments, we employ the ALERT dataset and RaDA dataset [6] which are open dataset leveraging IR-UWB to detect driver distraction activities.
@@
For model training, we strictly separate the training and testing datasets; the testing dataset consists of data from drivers who were not seen during the training phase. In addition, we employ a subject-wise leave-one-subject-out protocol: in each fold, all samples/segments from one driver are held out as the test set, and the remaining drivers are used for training. Thus, no driver appears in both training and testing, ensuring strict subject independence and preventing leakage across splits. We also note that we do not apply any data-dependent preprocessing such as normalization or standardization. We employ a learning rate as 0.00001 to ensure convergence the training loss. The base model is a ViT pre-trained on ImageNet [10], which uses embedding vectors of size 1,024. We fine-tune the model over 30 epochs, optimizing it with cross-entropy loss and the Adam optimizer to ensure robust performance.
@@
TABLE VI: Tester-wise standard deviation ($\sigma$) of learning algorithms.
Method GoogLeNet ResNet DenseNet MobileNet RNN DeiT ViT ISA-ViT
$\sigma$ 9.97 7.95 7.33 7.22 14.13 8.81 12.29 7.54
TABLE VII: Benchmarking [precision / recall] of learning algorithms
@@
VI-B Benchmarking Results for ALERT Dataset
VI-B1 Benchmarking of learning algorithms
For rigorous evaluation, we conduct benchmark tests using an unseen dataset during the training phase. In these experiments, we fine-tune pre-trained models, leveraging the pre-trained models’ great capabilities for feature extraction effectively to the UWB domain. This fine-tuning step is also essential to reduce the residual image-to-radar domain gap even after resolving the structural input-size and PEV mismatch.
To identify a suitable model for UWB DAR, we first present the average accuracy of each algorithm. We select four unseen testers and perform leave-one-out cross-validation to calculate their average accuracy. In each fold, the held-out tester is used only for testing, while all remaining testers are used for training, so that no subject overlaps between training and testing. Fig. 9(a) and Table. VI show a box plot and tester-wise standard deviation of various learning algorithms over three repeated runs (N=3) with different random seeds.
CNN-based approaches achieve 61.37–71.22% accuracy, outperforming both RNN- and transformer-based methods. The weaker performance of the RNN-based approach is mainly due to the absence of a pre-trained LSTM: although it uses pre-trained backbones for snippet-level feature extraction, the LSTM itself is trained from scratch, limiting overall effectiveness.
Transformer-based approaches also underperform CNNs in this UWB setting. While transformers are strong for image tasks, naïvely applying them to UWB requires resizing inputs and aligning pre-trained PEVs; without properly handling these issues (Sec. V), they cannot fully realize their potential.
In contrast, ISA-ViT addresses this mismatch by resizing UWB data without information loss and adapting pre-trained PEVs accordingly, achieving the best accuracy of 76.28%.
Beyond overall accuracy, we also report per-activity precision and recall, which provide complementary insight into classification performance. Table VII summarizes the precision and recall for each activity across the evaluated algorithms; the underlined entries highlight activities with notable fluctuations. Such variance is important because it indicates how consistently a model performs across different activities. Fig. 9(b) visualizes the mean and deviation of these metrics for each algorithm. As shown, several baselines exhibit large deviations, whereas ISA-ViT maintains relatively stable variance, resulting in more consistent per-activity performance.